Grade 12th passDifferential Calculus

compute the n-th differerntial coefficient of y=xloge^x

reese
9 Years agoGrade 12th pass

1 Answer

Nandana
9 Years ago
I think it is y= x log ex not y= x log e^x  !
      if that is  y= x log ex  then
             d / dx (y) = x (1/x) + ln x
                            = 1+ ln x
             d(2) / dx = 1/x
             d(3) / dx = -1/ x^2
             d(4) / dx = 2/ x^3
             .
             .
             .
             .
            d(n) / dxn =  (n-2)! / x^(n-1) , if  n is even if not (-) will come